# Cycle-Consistent Bridge Diffusion Model for Accelerated MRI Reconstruction

### v1

@@ 3.3 Contourlet Decomposition Embedding Module
The denoising networks of our CBDM model adopt a U-Net [31] with a time embedding. As illustrated in Fig. 3, we further propose a contourlet decomposition embedding module to enhance the backbone. Generally, the blurring and artifacts in under-sampled MR magnitude images are caused by the missing high-frequency information in the frequency space. Contourlet decomposition employs the Laplacian Pyramid (LP) [4] and Directional Filter Banks (DFB) [2, 7] to process the low-pass images iteratively. LP aims to achieve low- and high-frequency decomposition in a multi-scale fashion. DFB is used to reconstruct the original signal with minimal samples, producing minimal sample representations in the 2D frequency domain through j-level binary tree decomposition, resulting in $2^{j}$ directional sub-bands.
For example, the frequency domain is divided into $2^{3}$ directional sub-bands when $j=3$, and sub-bands $0-3$ and $4-7$ correspond to the vertical and horizontal details, respectively. Fig 4 illustrates the visualization of sub-band features at different levels obtained from the contourlet decomposition of $x_{t}$. It can be observed that contourlet decomposition effectively extracts high-frequency features of image edges. Then, the Contourlet Decomposition Embedding Module further resizes the features of multiple sub-bands at each level to match the feature size at the corresponding level, stacks them along the channel dimension, and then uses three convolution layers to align the channel number as corresponding level features.
@@
Figure 4: Contourlet decomposition features visualization with different level.
Figure 5: The reconstruction results (in-distribution) of fastMRI multi-coil knee data at equally-spaced Cartesian undersampling with an acceleration factor of 4 are shown. The first row displays the ground truth and the reconstructions from pGAN, cGAN, pix2pix, DDPM, IDDPM, AdaDiff, MC-DDPM, HFS-SDE and CBDM (Ours). The second row presents the enlarged view of the ROI, and the third row shows the error map of the reconstruction. The lower left corner contains the undersampling mask.
@@ 4.1 Data and Implementation Details
The experiments were conducted on the public benchmark fastMRI  [^1] http://fastmri.med.nyu.edu and IXI datasets  [^2] https://brain-development.org/ixi-dataset/. From the fastMRI dataset, we selected 360 individuals from the multi-coil knee dataset for training, 8 for validation, and 20 for testing. Following [5], the first six slices of each individual were excluded due to poor image quality. Consequently, the training, validation, and test sets contained 10,737, 244, and 576 2D images, respectively. Additionally, 218 images from 20 individuals in the fastMRI brain dataset were selected as an out-of-distribution set of images. All images were cropped to $320\times 320$ and normalized to the range [0, 1]. From the IXI dataset, we selected 577 patients with T1 images, randomly split into training (500 patients, 44,935 2D images), validation (37 patients, 3,330 images), and test (40 patients, 3,600 images) sets. All images were cropped to $256\times 256$ and normalized to the range [0, 1]. To simulate k-space undersampling, equally spaced Cartesian undersampling in the phase encoding direction was applied, with acceleration factors set to 4 and 8. The final MR magnitude images, both under-sampled and fully sampled, were generated from the corresponding k-space data through inverse Fourier transform and root sum of squares operations.
For training, the total number of iterations is 200k, using the AdamW optimizer with a learning rate of 1e-4 and a batch size of 28. For other diffusion methods (excluding AdaDiff), the training and sampling steps are set to 1000 and 200, respectively. In AdaDiff, the first stage employs an unconditional pretraining diffusion model with 1000 time steps, while the second stage uses 200 steps for iterative solving based on data consistency [46]. For our CBDM, both the training and sampling steps are set to 20, with the loss weight $\lambda$ set to 1. To ensure fairness, all diffusion models use the same network architecture. All experiments were conducted in the same environment using 4 NVIDIA Tesla V100 GPUs.
The comparison was conducted by the proposed CBDM model and pGAN [10], cGAN [48], pix2pix [21], DDPM [15], IDDPM [28], AdaDiff [14], MC-DDPM [39], HFS-SDE [5] and BBDM [25]. We further use the widely used Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM) and Normalized Mean Square Error (NMSE) for the measurement of reconstruction quality.
Figure 6: The reconstruction results of in-distribution IXI brain data at equally-spaced Cartesian undersampling with an acceleration factor of 4 are shown. The first row displays the ground truth and the reconstructions from pGAN, cGAN, pix2pix, DDPM, IDDPM, AdaDiff, MC-DDPM, HFS-SDE and CBDM (Ours). The second row presents the enlarged view of the ROI, and the third row shows the error map of the reconstruction. The lower left corner contains the undersampling mask.
@@
Figure 7: The reconstruction results (out-of-distribution) of fastMRI multi-coil brain data at equally-spaced Cartesian undersampling with an acceleration factor of 4 are shown. The first row displays the ground truth and the reconstructions from pGAN, cGAN, pix2pix, DDPM, IDDPM, AdaDiff, MC-DDPM, HFS-SDE and CBDM (Ours). The second row presents the enlarged view of the ROI, and the third row shows the error map of the reconstruction. The lower left corner contains the undersampling mask.
@@
Equi 4 IDDPM[28] 32.78±2.82 86.90±8.11 3.29±4.47
AdaDiff[14] 28.16±1.94 78.77±5.03 6.60±3.56
Diffusion MC-DDPM[39] 33.84±3.38 88.24±3.57 3.02±3.05
HFS-SDE[5] 30.70±2.10 82.19±5.36 3.79±3.24
BBDM[25] 33.79±2.48 87.85±7.19 2.47±3.04
CBDM(Ours) 34.84±2.48 89.73±6.20 1.96±2.40
pGAN[10] 27.23±1.82 71.34±8.47 7.55±5.28
GAN cGAN[48] 27.96±2.07 75.16±7.23 7.43±7.09
pix2pix[21] 30.80±2.00 82.44±6.95 4.29±4.50
@@
Equi 8 IDDPM[28] 30.84±2.44 81.87±8.92 4.60±5.40
AdaDiff[14] 28.85±2.15 77.53±6.15 6.04±4.58
Diffusion MC-DDPM[39] 31.30±2.45 84.58±5.84 3.30±2.79
HFS-SDE[5] 28.14±2.03 75.52±5.20 5.86±3.84
BBDM[25] 31.68±2.14 83.39±8.02 3.63±4.00
CBDM(Ours) 32.80±2.20 85.83±6.97 2.96±3.37
Table 1: Comparison of methods in in-distribution fastMRI knee experiments. Equi 4 and Equi 8 represent equally-spaced Cartesian undersampling with acceleration factors of 4 and 8, respectively. Bold and underline indicate the best and second-best results, respectively.
@@ 4.2 In-Distribution Results
Equi 4 IDDPM[28] 35.11±3.09 96.89±1.60 0.51±0.40
AdaDiff[14] 25.84±1.28 80.09±3.71 3.92±1.41
Diffusion MC-DDPM[39] 36.20±3.11 97.49±1.31 0.37±0.30
HFS-SDE[5] 30.76±2.80 89.22±2.15 1.16±0.60
BBDM[25] 36.57±3.10 97.70±1.25 0.36±0.29
CBDM(Ours) 37.56±3.07 98.10±1.02 0.29±0.22
pGAN[10] 24.15±5.28 89.75±1.84 5.01±1.33
GAN cGAN[48] 25.84±1.47 82.90±3.98 3.70±1.38
pix2pix[21] 31.12±1.96 93.94±2.01 1.12±0.55
@@
Equi 8 IDDPM[28] 34.89±3.02 96.53±1.73 0.53±0.40
AdaDiff[14] 23.74±1.05 75.45±3.87 6.45±1.86
Diffusion MC-DDPM[39] 35.53±2.96 97.04±1.53 0.45±0.34
HFS-SDE[5] 30.47±2.46 92.96±2.27 1.08±0.45
BBDM[25] 35.75±2.97 97.10±1.46 0.43±0.33
CBDM(Ours) 37.09±2.96 97.70±1.15 0.32±0.23
Table 2: Comparison of methods in in-distribution IXI experiments. Equi 4 and Equi 8 represent equally-spaced Cartesian undersampling with acceleration factors of 4 and 8, respectively. Bold and underline indicate the best and second-best results, respectively.
@@
Equi 4 IDDPM[28] 29.68±1.80 90.31±2.12 3.68±1.27
AdaDiff[14] 24.93±1.30 81.21±3.07 13.2±4.37
Diffusion MC-DDPM[39] 29.93±1.76 90.47±2.14 3.11±0.85
HFS-SDE[5] 26.03±1.35 79.69±3.58 7.40±1.34
BBDM[25] 29.41±1.64 91.01±1.78 3.16±1.12
CBDM(Ours) 30.67±1.71 91.97±1.42 2.98±0.94
pGAN[10] 24.15±1.06 71.10±5.19 12.5±3.27
GAN cGAN[48] 25.42±1.01 82.08±3.32 9.00±2.34
pix2pix[21] 25.72±1.13 79.32±4.71 9.41±2.69
@@
Equi 8 IDDPM[28] 27.36±1.30 86.95±2.19 6.35±1.64
AdaDiff[14] 25.67±1.37 80.08±3.13 10.0±2.72
Diffusion MC-DDPM[39] 26.10±2.21 82.52±3.56 7.91±2.89
HFS-SDE[5] 24.81±1.39 73.87±4.14 9.55±2.03
BBDM[25] 27.56±1.30 86.95±2.19 6.35±1.64
CBDM(Ours) 28.65±1.24 89.02±1.94 4.62±1.20
Table 3: Comparison of methods in out-of-distribution fastMRI brain experiments, using models trained on the in-distribution fastMRI knee dataset. Equi 4 and Equi 8 represent equally-spaced Cartesian under-sampling with acceleration factors of 4 and 8, respectively. Bold and underline indicate the best and second-best results, respectively.
Cycle-Consistent CDEM PSNR(dB)$\uparrow$ SSIM(%$\uparrow$ NMSE(%$\downarrow$
@@
Table 4: Ablation studies of different components. We report the metrics on fastMRI knee dataset under the Equi 4.
@@
$\sqrt{\sigma_{t1}}$ $\sqrt{\sigma_{t2}}$ PSNR(dB)$\uparrow$ SSIM(%$\uparrow$ NMSE(%$\downarrow$
1.0 1.0 34.67±2.50 89.54±6.20 2.03±2.49
1.0 1.1 34.79±2.54 89.73±6.19 1.97±2.41
1.0 1.2 34.84±2.54 89.73±6.20 1.96±2.40
1.0 1.3 34.76±2.56 89.70±6.19 2.00±2.47
Table 6: Ablation studies of different noise variances.
@@ 4.3 Out-of-Distribution Results
We further conducted out-of-distribution experiments to verify the generalization of our model. All models were trained on fastMRI knee data and tested by fastMRI brain MRI scans. The reconstruction results for equally-spaced Cartesian sampling with 4×4$\times$4 × acceleration are shown in Fig. 7. The reconstruction quality of pGAN, cGAN, and pix2pix methods significantly decreased, whereas diffusion-based methods maintained good reconstruction quality. Among them, our CBDM achieved the optimal reconstruction results with fewer noise and artifacts. Table 3 gives quantitative performance for the out-of-distribution experiments and our proposed CBDM model achieves the best performance for the brain data reconstruction.
@@ 4.4 Inference Efficiency
In the above experiments, the training and sampling steps for CBDM were set to 20 each, while other diffusion methods (except AdaDiff) used 1000 training steps and 200 sampling steps. The 200 sampling steps were accelerated using DDIM [34] sampling. For AdaDiff, the first stage involved pretraining with 1000 steps, and the second stage did not use 200 sampling steps but instead performed 200 online iterations based on the data consistency loss. As shown in Fig. 8 (a), CBDM’s significantly reduced sampling steps result in much higher inference efficiency compared to other diffusion methods.
To further compare inference efficiency fairly, we also set the sampling steps or the second-stage iteration steps of other methods to 20, as shown in Fig. 8 (b). This adjustment led to a substantial drop in reconstruction performance for these methods. Although using DDIM [34] with very few sampling steps can improve sampling speed, it comes with significant performance degradation, making it impractical for real-world applications. In contrast, CBDM achieves both training and sampling with only 20 steps, maintaining high efficiency without compromising reconstruction performance.
Figure 8: (a) A bar chart comparing the inference time of various diffusion methods. (b) A bar chart showing the changes in PSNR performance of other SOTA methods when their sampling steps are adjusted to match those of CBDM.

### v2

@@ 3.3 Contourlet Decomposition Embedding Module
As a signal representation tool, contourlets (analogous to edges in images) can effectively exploit geometric regularities to achieve sparse representations. Previous studies have utilized Fast Fourier Transform (FFT) or wavelet decomposition to capture texture information. However, the effectiveness of these methods is often limited in practice due to the lack of more precise design. As illustrated in Fig. 4(a), wavelet decomposition produces a large number of inefficient point-based decompositions. In contrast, Fig. 4(b) demonstrates that contourlet decomposition is capable of representing smooth contours using fewer coefficients by approximating them with linear segments. Consequently, contourlet-based representations are sparser and offer the advantages of directionality, locality, and bandpass characteristics.
@@
The denoising networks of our SC-NDB model adopt a U-Net [33] with a time embedding. As illustrated in Fig. 5, we further propose a contourlet decomposition embedding module to enhance the backbone. Generally, the blurring and artifacts in under-sampled MR magnitude images are caused by the missing high-frequency information in the frequency space. Contourlet decomposition employs the Laplacian Pyramid (LP) [5] and Directional Filter Banks (DFB) [2, 9] to process the low-pass images iteratively. The LP module performs a multi-scale decomposition to separate low- and high-frequency components, as illustrated in Fig. 6. Specifically, the low-pass subband is generated by applying a low-pass analysis filter to the input $x_{t}$, followed by downsampling via a sampling matrix $\mathbf{S}$. The high-pass subband is then derived as the residual between the original input $x_{t}$ and the reconstructed low-pass subband, which is obtained through upsampling and a low-pass synthesis filter. To further enhance the representation of directional features, DFB is employed to decompose the high-pass subband into multiple directional components. By utilizing a $j$-level binary tree structure, DFB produces $2^{j}$ directional subbands, yielding compact and directionally sensitive representations in the 2D frequency domain.
For example, the frequency domain is divided into $2^{3}$ directional sub-bands when $j=3$, and sub-bands $0-3$ and $4-7$ correspond to the vertical and horizontal details, respectively.
$F_{l},F_{h}$ $=LP(F_{l})\downarrow d$ (12)
$F_{bds}$ $=DFB(F_{h})\;n\in[1,j],$
where symbol $\downarrow$ is the downsampling operator and $d$ denotes the interlaced downsampling factor. The subscripts $l$ and $h$ represent the low-pass and high-pass components respectively, $bds$ indicates the bandpass directional subbands.
Figure 6: Details of the Laplacian Pyramid (LP) decomposition.
Fig 7 illustrates the visualization of sub-band features at different levels obtained from the contourlet decomposition of $x_{t}$. It can be observed that contourlet decomposition effectively extracts high-frequency features of image edges. Then, the Contourlet Decomposition Embedding Module further resizes the features of multiple sub-bands at each level to match the feature size at the corresponding level, stacks them along the channel dimension, and then uses three convolution layers to align the channel number as corresponding level features.
Figure 7: Contourlet decomposition features visualization with different levels.
@@ 4.1 Data and Implementation Details
The experiments were conducted on the public benchmark fastMRI  [^1] http://fastmri.med.nyu.edu and IXI datasets  [^2] https://brain-development.org/ixi-dataset/. On the fastMRI dataset, we selected 360 individuals from the multi-coil knee dataset for training, 8 for validation, and 20 for testing. Following [6], the first six slices of each individual were excluded due to poor image quality. Consequently, the training, validation, and test sets contained 10,737, 244, and 576 2D images, respectively. Additionally, 218 images from 20 individuals from the fastMRI brain dataset were selected as an out-of-distribution set of images. All images were cropped to $320\times 320$ and normalized to the range [0, 1]. From the IXI dataset, we selected 577 patients with T1 images, randomly split into training (500 patients, 44,935 2D images), validation (37 patients, 3,330 images), and test (40 patients, 3,600 images) sets. All images were cropped to $256\times 256$ and normalized to the range [0, 1]. To simulate k-space undersampling, equally spaced Cartesian undersampling in the phase encoding direction was applied, with acceleration factors set to 4 and 8. The final MR magnitude images, both under-sampled and fully sampled, were generated from the corresponding k-space data through inverse Fourier transform and root sum of squares operations.
@@
Figure 8: The reconstruction results of magnitude images generated from multi-coil knee fast MRI with a 4× acceleration factor under uniform Cartesian undersampling are presented (in-distribution). The first row displays the ground truth and the reconstructions from DDPM, I2SB, BBDM, selfRDB and SC-NDB (Ours). The second row presents the enlarged view of the ROI, and the third row shows the error map of the reconstruction. The lower left corner contains the undersampling mask.
For training, the total number of iterations is 200k, using the AdamW optimizer with a learning rate of 1e-4 and a batch size of 28. For other diffusion methods (excluding AdaDiff), the training and sampling steps are set to 1000 and 200, respectively. In AdaDiff, the first stage employs an unconditional pretraining diffusion model with 1000 time steps, while the second stage uses 200 steps for iterative solving based on data consistency [46]. For our SC-NDB, both the training and sampling steps are set to 20, with the loss weight $\lambda$ set to 1. All experiments were conducted in the same environment using 4 NVIDIA Tesla V100 GPUs.
The comparison was conducted by the proposed SC-NDB model and AdaDiff [16], MC-DDPM [40], HFS-SDE [6], DDPM [17], BBDM [25], I2SB [26], and selfRDB [1], where AdaDiff, MC-DDPM, and HFS-SDE are non-magnitude-image-based approaches. We further use the widely used Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM) and Normalized Mean Square Error (NMSE) for the measurement of reconstruction quality.
Figure 9: The reconstruction results of magnitude images generated from multi-coil brain fast MRI with a 4× acceleration factor under uniform Cartesian undersampling are presented (out-of-distribution). The first row displays the ground truth and the reconstructions from DDPM, I2SB, BBDM, selfRDB and SC-NDB (Ours). The second row presents the enlarged view of the ROI, and the third row shows the error map of the reconstruction. The lower left corner contains the undersampling mask.
@@
4×4$\times$4 × I2SB[26] 33.45±2.54 87.49±6.53 2.59±3.08
Magnitude BBDM[25] 33.79±2.48 87.85±7.19 2.47±3.04
selfRDB[1] 34.20±2.50 89.19±6.28 2.26±2.80
SC-NDB(Ours) 34.84±2.48 89.73±6.20 1.96±2.40
AdaDiff[16] 28.85±2.15 77.53±6.15 6.04±4.58
Non-magnitude MC-DDPM[40] 31.30±2.45 84.58±5.84 3.30±2.79
HFS-SDE[6] 30.88±2.47 83.87±5.92 3.60±5.40
@@
8×8$\times$8 × I2SB[26] 31.45±2.25 82.54±7.73 3.78±4.03
Magnitude BBDM[25] 31.68±2.14 83.39±8.02 3.63±4.00
selfRDB[1] 31.98±2.18 84.88±7.17 3.41±3.76
SC-NDB(Ours) 32.80±2.20 85.83±6.97 2.96±3.37
Table 1: Comparison of methods in in-distribution fastMRI knee experiments. 4×4$\times$4 × and 8×8$\times$8 × represent equally-spaced Cartesian undersampling with acceleration factors of 4 and 8, respectively. Bold and underline indicate the best and second-best results, respectively.
@@ 4.2 In-Distribution Results
Fig. 8 illustrates representative reconstruction results from the fastMRI knee dataset under 4×4$\times$4 × equally spaced Cartesian sampling. The first row displays outputs from DDPM, I2SB, BBDM, selfRDB, and SC-NDB. The second row presents enlarged views of selected regions of interest (ROI), while the third row shows the corresponding error heatmaps. Reconstructions generated by DDPM, I2SB, BBDM, and selfRDB exhibit noticeable artifacts and noise, often leading to the loss of fine anatomical structures. In contrast, SC-NDB demonstrates superior reconstruction fidelity by preserving realistic high-frequency details and effectively suppressing noise and artifacts.
@@
4×4$\times$4 × I2SB[26] 34.31±2.76 96.23±1.71 0.58±0.41
Magnitude BBDM[25] 36.57±3.10 97.70±1.25 0.36±0.29
selfRDB[1] 36.02±3.02 97.49±1.31 0.41±0.31
SC-NDB(Ours) 37.56±3.07 98.10±1.02 0.29±0.22
AdaDiff[16] 23.74±1.05 75.45±3.87 6.45±1.86
Non-magnitude MC-DDPM[40] 35.53±2.96 97.04±1.53 0.45±0.34
HFS-SDE[6] 34.99±3.04 96.58±1.77 0.52±0.41
@@
8×8$\times$8 × I2SB[26] 33.11±2.30 95.30±1.74 0.73±0.41
Magnitude BBDM[25] 35.75±2.97 97.10±1.46 0.43±0.33
selfRDB[1] 35.38±2.93 97.01±1.49 0.47±0.35
SC-NDB(Ours) 37.09±2.96 97.70±1.15 0.32±0.23
Table 2: Comparison of methods in in-distribution IXI experiments. 4×4$\times$4 × and 8×8$\times$8 × represent equally-spaced Cartesian undersampling with acceleration factors of 4 and 8, respectively. Bold and underline indicate the best and second-best results, respectively.
@@
4×4$\times$4 × I2SB[26] 28.12±1.97 86.98±3.41 4.83±2.10
Magnitude BBDM[25] 29.41±1.64 91.01±1.78 3.16±1.12
selfRDB[1] 29.78±1.64 90.95±2.08 3.49±1.25
SC-NDB(Ours) 30.67±1.71 91.97±1.42 2.98±0.94
AdaDiff[16] 25.67±1.37 80.08±3.13 10.0±2.72
Non-magnitude MC-DDPM[40] 26.10±2.21 82.52±3.56 7.91±2.89
HFS-SDE[6] 26.36±2.10 83.35±2.32 7.35±1.62
@@
8×8$\times$8 × I2SB[26] 25.77±1.30 82.48±4.00 7.67±2.16
Magnitude BBDM[25] 27.56±1.30 86.95±2.19 6.35±1.64
selfRDB[1] 27.46±1.24 87.46±2.66 6.04±1.71
SC-NDB(Ours) 28.65±1.24 89.02±1.94 4.62±1.20
Table 3: Comparison of methods in out-of-distribution fastMRI brain experiments, using models trained on the in-distribution fastMRI knee dataset. 4×4$\times$4 × and 8×8$\times$8 × represent equally-spaced Cartesian under-sampling with acceleration factors of 4 and 8, respectively. Bold and underline indicate the best and second-best results, respectively.
Figure 10: The Wilcoxon signed-rank test with a significance level of $p<0.005$ was conducted to assess the statistical differences between SC-NDB and both BBDM and SelfRDB.
Self-Consistency CDEM PSNR(dB)$\uparrow$ SSIM(%$\uparrow$ NMSE(%$\downarrow$
@@
Table 4: Ablation studies of different components. We report the metrics on the fastMRI knee dataset at a 4× acceleration factor.
@@
$\sqrt{\sigma_{t2}}$ PSNR(dB)$\uparrow$ SSIM(%$\uparrow$ NMSE(%$\downarrow$
1.0 34.67±2.50 89.54±6.20 2.03±2.49
1.1 34.79±2.54 89.73±6.19 1.97±2.41
1.2 34.84±2.54 89.73±6.20 1.96±2.40
1.3 34.76±2.56 89.70±6.19 2.00±2.47
Table 6: Ablation study was performed by varying the noise variance using $\sqrt{\sigma_{t2}}$ to evaluate its impact on reconstruction performance.
@@ 4.3 Out-of-Distribution Results
We further conducted out-of-distribution experiments to verify the generalization of our model. All models were trained on fastMRI knee data and tested by fastMRI brain MRI scans. The reconstruction results for equally-spaced Cartesian sampling with 4×4$\times$4 × acceleration are shown in Fig. 9. The reconstruction quality of DDPM, I2SB, BBDM and selfRDB methods significantly decreased, whereas diffusion-based methods maintained good reconstruction quality. Among them, our SC-NDB achieved the optimal reconstruction results with fewer noise and artifacts. Table 3 gives quantitative performance for the out-of-distribution experiments and our proposed SC-NDB model achieves the best performance for the brain data reconstruction.
@@ 4.4 Inference Efficiency
In the above experiments, the training and sampling steps for SC-NDB and selfRDB were set to 20 each, while other diffusion methods (except AdaDiff) used 1000 training steps and 200 sampling steps. The 200 sampling steps were accelerated using DDIM [36] sampling. For AdaDiff, the first stage involved pretraining with 1000 steps, and the second stage did not use 200 sampling steps but instead performed 200 online iterations based on the data consistency loss. As shown in Fig. 11, SC-NDB’s significantly reduced sampling steps result in much higher inference efficiency compared to other diffusion methods.
Figure 11: A bar chart comparing the inference time of various diffusion methods.